Science News
MOS Mask Milling Machine Issue – Resolved
Posted Apr 10, 2025
April 11, 2025: Problem with the oxygen regulator was found and resolved. The MOS mask milling machine has been returned to operation.
April 9, 2025: The laser for the MOS mask milling machine recently stopped working. Currently all mask milling for LUCI and MODS is halted while the issue is resolved.
LUCI1 Filter Wheel#2 (Broadband) Switch Issue
Posted Apr 03, 2025
Apr 2, 2025: Flaky switch on LUCI1 Filter Wheel #2 (broadband filters).
LUCI1’s Filter Wheel #2, where the broadband filters are located, is showing a flaky position sensing switch that affects all higher-numbered filters, which includes all but the zJspec and J filters. LBTO assistance will be needed to safely drive LUCI observations until this gets repaired.
LUCI1 Critical MOS error Recovery Status
Posted Mar 09, 2025
Mar 13, 2025: All normal modes with LUCI1 are released for science.
N375 and N30 imaging, as well as G210 and G200 spectroscopic modes have all been released for normal science use.
Mar 9, 2025: LUCI1 All modes **NOT** using G210 or G040 are released for science.
LUCI recalibration is underway. All modes **NOT** using G210 or G040 are released for science.
LUCI1 is working well mechanically and all the major modes have new flexure tables. In G210 there is still an issue with the wavelength calibration which is being investigated and reviewed.
Feb 13, 2025: LUCI1 Troubleshooting/Cooling in the lab
Maintenance work and testing continued through the month of January.and beginning of February. The MOS translator issue was addressed, as well as issues with the LUCI1 grating turret. Other general maintenance items were also taken care of while the instrument was disassembled. LUCI1 has passed a full warm checkout and is now cooling in the lab.
Jan 06, 2025: LUCI1 Lab Troubleshooting
LUCI1 troubleshooting in the lab is ongoing. The MOS unit is now able to reliable move the mask with the unit warm. Several motors are scheduled for replacement.
Dec 30, 2024: LUCI1 Removed from telescope
The MOS troubleshooting revealed an issue internal to the instrument. LUCI1 was removed from the telescope to allow for continued troubleshooting and repairs in the lab.
Dec 9, 2024: LUCI1 Offline with MOS error
LUCI1 experience a critical MOS error on December 09, 2024. LUCI1 taken offline for troubleshooting.
SX ASM Status Update
Posted Feb 14, 2025
Update Feb 24, 2025
On-sky seeing limited checkouts were completed on UTFeb24. The binodal astigmatism was measured and collimation tables updated. The SX ASM is released for all available seeing-limited science modes without restrictions.
Update Feb 14, 2025
ASM and AO calibration and checkout teams were able to swiftly integrate, calibrate, and checkout the unit for SHARK-NIR science. The SX AO has been released for shared-risk SHARK-NIR science.
LUCI1 remains off the telescope (see here for details). The AO team is working to test the SX LUCI WFS and SX ASM with IRTC1 in place of LUCI1 to ensure the system is ready when LUCI1 becomes operational.
Modes available for shared-risk release include:
- SX PEPSI PFU
- MODS1 Longslit Spectroscopy
- SHARK-NIR AO
Wide field imaging and MOS spectroscopy remain unavailable pending the measurement of the field aberration.
Update Feb 08, 2025:
The TS5 was partially realuminized to recoat the impacted area on Jan 25, 2025 and reintegrated with the SX ASM on Jan 31, 2025. After passing lab tests, the unit was installed on the telescope on Feb 07, 2025. The flat was verified for seeing limited operations and the unit was released for seeing-limited shared-risk science.
The science modes available for shared risk seeing limited science include SX PEPSI PFU, and MODS1 for longslit spectroscopy with the caveat that sources may not be optimally collimated. MODS imaging and MODS MOS Spectroscopy should wait until we can execute the field aberration work. LUCI1 is pending installation on the telescope (see note here for details).
Update Jan 09, 2025:
Several actuators were replaced on the SX ASM. TS5 recoating operation is scheduled for the week of January 20 at Sunnyside. The ASM team is preparing SX ASM to integrate with the recoated TS5 the week of January 27.
Update Sept 20, 2024:
Assessments by the AO team have revealed that the rear-side surface of the SX thin shell (TS5) suffered significant water damage that will very likely require recoating. Planning for recoating the thin shell is happening now and the optimistic timeline is that this would enable the SX ASM to return to service sometime in November. The AO team has now shifted efforts to the DX ASM.
Update Sept 9, 2024:
The preliminary inspection by the ASM team shows signs of watermarks and possible burns in one of the distribution boards. After disconnecting the affected board, the team powered on the rest of the unit and measured the gap for all but ~ 6% of the thin shell’s area. The measurements suggest that the water intrusion is receding (leaving the gap), which is encouraging. The team will thoroughly assess the water damage when they lower the thin shell, which is expected around September 17.
Initial Note: Sept 2, 2024
On UT Sept 1 the telescope experienced a water intrusion event. The SX ASM thin shell had liquid gap contamination impacting the outer rings between 7-9 O’clock. The unit was removed on Sept 2 for repairs. The timeline for repairs is TBD.
The rigid secondary remains installed in place of the DX side while the DX ASM campaign wraps up. The SX side of the telescope will remain without a secondary mirror until the DX ASM is installed later this month.
DX AO in Cold Conditions
Posted Jan 21, 2025
DX ASM in Cold Weather Update – Jan 21, 2025
Work was done in summer shutdown 2024 to attempt to address the issues causing the failures in cold temperatures with the DX ASM. Unfortunately, we have continued to see the failures as we resume operations in cold temperatures.
If the ICS (Instrument cooling System) temperature drops below -7 Celsius (nominally chamber temperature dropping below -5 Celsius, unless the ICS offset is adjusted), the AO operation should be suspended until the ICS temperature rises above -5 Celsius (nominally chamber temperature rising above -3 Celsius)
Changes to the instrument cooling setpoint can impact seeing limited and other instrument operations and is done with care. Users with AO programs are encouraged to coordinate their scheduled AO observations with LBTO.
DX in Cold Weather – Update (Jan 6 2024)
The failure is linked to the cooling temperature at the adaptive secondary. The cooling setpoint was slightly adjusted to provide additional leeway while maintaining the safety of the adsec and neighboring systems.
For the safety of the system, DX AO operations should be halted if the chamber temperatures fall below -6C.
Nov 27, 2023 – DX in Cold Weather
Last February we experienced a number of failures of the DX AO at cold temperatures. At temperatures below -5C, there has been a high frequency of communication failures. The main signature of these failures is that the FastDiagnostics data stop being received by the DX ASM control computer. The result of this interruption is that the control computer forces a DISABLE_COIL which rests the shell and stops DX AO operation. The issue causing this particular failure remains unresolved at this time, and for the safety of the system operational restrictions have been put in place.
For the safety of the system, DX AO operations should be halted if the chamber temperatures fall below -5C.
The team continues to review data to investigate the precise temperature threshold of these events, as well as if other environmental or operational factors (such as loop frequency) may play a role. An off-telescope intervention will likely be required to fully resolve this.
DX LUCI AO FW issue – Shared Risk Release
Posted Jan 11, 2025
Jan 09, 2025
The DX LUCI WFS experienced an issue with one of the filter wheels. This filter wheel is used for very bright targets only with an R magnitude brighter than 8th magnitude. LUCI full AO and LUCI ESM is released in shared risk. PI’s with bright targets are encouraged to reach out to the science operations team to verify configurations.